OpenStudy (anonymous):

Solve for r: r − 12 = 11

OpenStudy (konradzuse):

r - 12 = 11 r = 11+12 r = 23

OpenStudy (anonymous):

KonradZuse is correct!!

OpenStudy (konradzuse):

I don't mind helping you, but try to answer them on your own as well, to learn :)

OpenStudy (hang254):

^^true

OpenStudy (hang254):

not that, konrads answer